When a film of transparent material of refractive index1.2 is put behind one of the slits of a two-slit Young's experiment,the zeroth order bright fringe moves to the position previously occupied by the 4 the order bright fringe.Given that the wavelength of the light used 5000 Angstrom.Find the thickness of the film.
 

Grade:12th pass

2 Answers

Satya
4 years ago
Extra path difference for Central position =(ū-1)t
As it coincides with previous 4th bright
(ū-1)t = 4 lambda
t = 10-5 m
 
Arun
3 years ago
Dear student
 
 
Extra path difference for Central position =(ū-1)t
As it coincides with previous 4th bright
(ū-1)t = 4 lambda
t = 10-5 m
 
Hope it helps